# Best Arm Identification for Prompt Learning under a Limited Budget

### v2

@@ 5.1 Evaluating TRIPLE with Fixed Prompt Pools
Cause and effect 0.65 $\pm$ 0.18 0.74 $\pm$ 0.06 0.75 $\pm$ 0.13 0.78 $\pm$ 0.08 0.78 $\pm$ 0.15 0.80 $\pm$ 0.05 0.80 $\pm$ 0.08
Common concept 0.09 $\pm$ 0.05 0.12 $\pm$ 0.06 0.10 $\pm$ 0.04 0.14 $\pm$ 0.05 0.12 $\pm$ 0.04 0.12 $\pm$ 0.05 0.14 $\pm$ 0.01
Disambiguation qa 0.83 $\pm$ 0.04 0.88 $\pm$ 0.10 0.97 $\pm$ 0.01 0.96 $\pm$ 0.01 0.95 $\pm$ 0.04 0.98 $\pm$ 0.02 0.96 $\pm$ 0.02
Gender inc. DE 0.74 $\pm$ 0.17 0.81 $\pm$ 0.10 0.85 $\pm$ 0.12 0.84 $\pm$ 0.14 0.69 $\pm$ 0.22 0.80 $\pm$ 0.17 0.88 $\pm$ 0.05
Hyperbaton 0.78 $\pm$ 0.07 0.83 $\pm$ 0.11 0.84 $\pm$ 0.12 0.84 $\pm$ 0.11 0.59 $\pm$ 0.24 0.74 $\pm$ 0.21 0.79 $\pm$ 0.18
Larger animal 0.56 $\pm$ 0.24 0.64 $\pm$ 0.25 0.79 $\pm$ 0.06 0.84 $\pm$ 0.02 0.66 $\pm$ 0.13 0.73 $\pm$ 0.18 0.85 $\pm$ 0.15
Movie recommendation 0.61 $\pm$ 0.12 0.65 $\pm$ 0.18 0.76 $\pm$ 0.06 0.74 $\pm$ 0.14 0.67 $\pm$ 0.11 0.65 $\pm$ 0.15 0.71 $\pm$ 0.15
Object counting 0.41 $\pm$ 0.12 0.45 $\pm$ 0.08 0.50 $\pm$ 0.07 0.48 $\pm$ 0.12 0.44 $\pm$ 0.08 0.50 $\pm$ 0.09 0.49 $\pm$ 0.07
Orthography starts with 0.41 $\pm$ 0.21 0.65 $\pm$ 0.16 0.67 $\pm$ 0.12 0.66 $\pm$ 0.13 0.58 $\pm$ 0.13 0.64 $\pm$ 0.09 0.67 $\pm$ 0.17
Question selection 0.90 $\pm$ 0.04 0.91 $\pm$ 0.03 0.95 $\pm$ 0.01 0.93 $\pm$ 0.03 0.93 $\pm$ 0.06 0.92 $\pm$ 0.06 0.93 $\pm$ 0.03
Rhymes 0.66 $\pm$ 0.30 0.68 $\pm$ 0.26 0.75 $\pm$ 0.20 0.78 $\pm$ 0.16 0.78 $\pm$ 0.12 0.83 $\pm$ 0.08 0.85 $\pm$ 0.13
Snarks 0.44 $\pm$ 0.10 0.52 $\pm$ 0.19 0.57 $\pm$ 0.10 0.60 $\pm$ 0.21 0.49 $\pm$ 0.17 0.56 $\pm$ 0.15 0.67 $\pm$ 0.05
@@ 5.2 Integrating TRIPLE into End-to-End Pipelines
We now explore whether TRIPLE can provide performance improvements when plugged into end-to-end prompt learning pipelines that include both prompt generation and selection. To this end, two end-to-end designs are considered, aiming to assess the performance of TRIPLE in more fluid and iterative settings, which are discussed in the following with our implementation details.
  • APE. Proposed by Zhou et al., (2022), the APE pipeline lets LLMs generate prompt candidates and then selects from them. In our experiments, for each task, following original templates, $30$ prompts are generated, followed by different methods to perform selection with a budget of $5$ evaluations per prompt on average (i.e., $150$ LLM accesses overall). The authors of Zhou et al., (2022) suggest a non-iterative version with uniform evaluations of prompts, which is taken as the baseline here.
@@
The end-to-end experiment results are reported in Table 2, which reveal the consistently better performance of TRIPLE over the originally adopted baseline methods. This observation highlights the applicability and flexibility of the TRIPLE framework, i.e., it can benefit any prompt learning pipelines requiring a selection component.
6 Related Works
Prompt Learning. The study of prompt learning (also known as instruction learning) focuses on automatically learning suitable prompts, which is more scalable compared with manual prompt engineering. Many efforts have been devoted to this direction (Gao et al.,, 2020; Wen et al.,, 2023; Sun et al.,, 2023; Zhang et al.,, 2022; Zhang et al., 2023b, ), which is summarized in a recent survey of Liu et al., (2023). The early studies mostly consider optimizing soft prompts (i.e., continuous vectors) (Lester et al.,, 2021; Li and Liang,, 2021; Zhong et al.,, 2021; Liu et al.,, 2021; Wu et al.,, 2022) or discrete but not interpretable prompts (Shin et al.,, 2020; Shi et al.,, 2022) for white-box LLMs, where different gradient-guided optimization techniques are leveraged. The recent research efforts, instead, are more focused on considering the more practical setting of learning interpretable prompts for black-box LLMs (Prasad et al.,, 2022; Zhou et al.,, 2022; Pryzant et al.,, 2023; Xu et al.,, 2022; Guo et al.,, 2023; Pan et al.,, 2023; Chen et al.,, 2023), where the generating-then-selecting pipeline in Fig. 1 is often adopted.
Compared with previous investigations, this work additionally considers a limited budget, which we believe is a practical but under-investigated concern. Most of the previous methods perform selection based on evaluating prompts on all development data (Jiang et al.,, 2020; Xu et al.,, 2022; Guo et al.,, 2023; Prasad et al.,, 2022), which is unavoidably costly. APE (Zhou et al.,, 2022) briefly touched on the cost issue by proposing a naive iterative top filtering strategy (which however is suggested to have only marginal benefits). APO (Pryzant et al.,, 2023) tested a few MAB methods, including two classical BAI-FB designs, i.e., SH (Karnin et al.,, 2013) and SR (Audibert et al.,, 2009), and reported that the performance of UCB is the more favorable. However, it neither formally introduces the budget constraint nor provides a systematical connection to BAI-FB as in this work. Also, this work goes much deeper in incorporating the state-of-the-art BAI-FB designs (i.e., CR (Wang et al.,, 2023) and GSE (Azizi et al.,, 2023)) and proposing embedding-based enhancements. It is worth noting that INSTINCT (Lin et al.,, 2023) also incorporates one MAB method, i.e., NeuralUCB (Zhou et al.,, 2020), to prompt learning. However, Lin et al., (2023) requires a white-box LLM while NeuralUCB is designed for regret minimization (instead of best arm identification), which is not suitable for learning the optimal prompt.
Multi-armed Bandits. Here we briefly discuss the representative studies of MAB, with comprehensive surveys available in Bubeck et al., (2012); Lattimore and Szepesvári, (2020). As mentioned in Sec. 3.1, the target of learning in a MAB system can be roughly categorized as regret minimization and best arm identification. The regret minimization designs target achieving a desired balance between exploration and exploitation so that the cumulative rewards are maximized, e.g., UCB (Auer et al.,, 2002) and Thompson sampling (Thompson,, 1933). The best arm identification (BAI) designs are fully focused on exploration and can be further divided into two classes: fixed-budget (BAI-FB) and fixed-confidence (BAI-FC). The BAI-FB setting maximizes the probability of finding the best arm with a limited number of pulls (Audibert et al.,, 2010; Karnin et al.,, 2013; Wang et al.,, 2023; Azizi et al.,, 2023; Atsidakou et al.,, 2022; Yang and Tan,, 2022). The BAI-FC setting is a dual one which focuses on minimizing the overall number of pulls while guaranteeing that the probability of finding the best arm is higher than a threshold (Garivier and Kaufmann,, 2016; Jamieson and Nowak,, 2014; Soare et al.,, 2014; Jedra and Proutiere,, 2020). Besides leveraging BAI-FB as in this work, it is imaginable that BAI-FC designs can also find applications in prompt learning, especially when valuing identification accuracy over incurred costs. While MAB has found wide success in different applications, this work marks the first time that a systematical connection between MAB and prompt learning has been established to the best of our knowledge.
@@ 7 Conclusions
Prompt learning is an important problem for large language models (LLMs), but prior research has not considered the potential cost during prompt selection. We have explicitly incorporated a budget constraint to prompt learning, and studied the problem of how to efficiently select prompts with the given budget. A systematical connection between multi-armed bandits (MAB) and prompt learning was established. Through this lens, we proposed a general framework, termed TRIPLE, to fully harness the power of fixed-budget best arm identification (BAI-FB) to perform prompt learning. Besides standard BAI-FB designs, two embedding-based enhancements were proposed to accelerate learning when selecting from large prompt pools. Extensive experimental results demonstrated the superiority of TRIPLE on prompt selection in multiple representative tasks and various targeted LLMs. Furthermore, we showed that TRIPLE could be plugged into popular end-to-end prompt learning pipelines, with better performance than previous implementations, demonstrating its effectiveness and flexibility.
@@
In addition to the technical contributions, we believe that the connection between prompt learning and MAB may be of broader interest. It not only provides a rich set of tools from MAB to advance prompt learning but also introduces a new application scenario for MAB (especially best arm identification) research. For example, the research on contextual bandits (Li et al.,, 2010) may provide insights into selecting input-dependent prompts (Wu et al.,, 2022). Also, the new application of prompt learning may spark new research efforts in MAB, e.g., developing more efficient BAI methods for correlated arms.
Appendix A Details of the Adopted BAI-FB Designs

### v3

@@ 5.1 Evaluating TRIPLE with Fixed Prompt Pools
(#1) Cause and effect 0.65 $\pm$ 0.18 0.74 $\pm$ 0.06 0.75 $\pm$ 0.13 0.78 $\pm$ 0.08 0.78 $\pm$ 0.15 0.80 $\pm$ 0.05 0.80 $\pm$ 0.08
(#2) Common concept 0.09 $\pm$ 0.05 0.12 $\pm$ 0.06 0.10 $\pm$ 0.04 0.14 $\pm$ 0.05 0.12 $\pm$ 0.04 0.12 $\pm$ 0.05 0.14 $\pm$ 0.01
(#3) Disambiguation qa 0.83 $\pm$ 0.04 0.88 $\pm$ 0.10 0.97 $\pm$ 0.01 0.96 $\pm$ 0.01 0.95 $\pm$ 0.04 0.98 $\pm$ 0.02 0.96 $\pm$ 0.02
(#4) Gender inc. DE 0.74 $\pm$ 0.17 0.81 $\pm$ 0.10 0.85 $\pm$ 0.12 0.84 $\pm$ 0.14 0.69 $\pm$ 0.22 0.80 $\pm$ 0.17 0.88 $\pm$ 0.05
(#5) Hyperbaton 0.78 $\pm$ 0.07 0.83 $\pm$ 0.11 0.84 $\pm$ 0.12 0.84 $\pm$ 0.11 0.59 $\pm$ 0.24 0.74 $\pm$ 0.21 0.79 $\pm$ 0.18
(#6) Larger animal 0.56 $\pm$ 0.24 0.64 $\pm$ 0.25 0.79 $\pm$ 0.06 0.84 $\pm$ 0.02 0.66 $\pm$ 0.13 0.73 $\pm$ 0.18 0.85 $\pm$ 0.15
(#7) Movie recommendation 0.61 $\pm$ 0.12 0.65 $\pm$ 0.18 0.76 $\pm$ 0.06 0.74 $\pm$ 0.14 0.67 $\pm$ 0.11 0.65 $\pm$ 0.15 0.71 $\pm$ 0.15
(#8) Object counting 0.41 $\pm$ 0.12 0.45 $\pm$ 0.08 0.50 $\pm$ 0.07 0.48 $\pm$ 0.12 0.44 $\pm$ 0.08 0.50 $\pm$ 0.09 0.49 $\pm$ 0.07
(#9) Orthography starts with 0.41 $\pm$ 0.21 0.65 $\pm$ 0.16 0.67 $\pm$ 0.12 0.66 $\pm$ 0.13 0.58 $\pm$ 0.13 0.64 $\pm$ 0.09 0.67 $\pm$ 0.17
(#10) Question selection 0.90 $\pm$ 0.04 0.91 $\pm$ 0.03 0.95 $\pm$ 0.01 0.93 $\pm$ 0.03 0.93 $\pm$ 0.06 0.92 $\pm$ 0.06 0.93 $\pm$ 0.03
(#11) Rhymes 0.66 $\pm$ 0.30 0.68 $\pm$ 0.26 0.75 $\pm$ 0.20 0.78 $\pm$ 0.16 0.78 $\pm$ 0.12 0.83 $\pm$ 0.08 0.85 $\pm$ 0.13
(#12) Snarks 0.44 $\pm$ 0.10 0.52 $\pm$ 0.19 0.57 $\pm$ 0.10 0.60 $\pm$ 0.21 0.49 $\pm$ 0.17 0.56 $\pm$ 0.15 0.67 $\pm$ 0.05
Avg. Performance Rank 4.00 $\pm$ 0.00 2.92 $\pm$ 0.28 1.58 $\pm$ 0.64 1.50 $\pm$ 0.50 2.75 $\pm$ 0.43 2.00 $\pm$ 0.71 1.25 $\pm$ 0.43
@@ 5.2 Integrating TRIPLE into End-to-End Pipelines
We now explore whether TRIPLE can provide performance improvements when plugged into end-to-end prompt optimization pipelines that include both prompt generation and selection. To this end, two end-to-end designs are considered, aiming to assess the performance of TRIPLE in more fluid and iterative settings, which are discussed in the following with our implementation details.
  • APE. Proposed by Zhou et al., (2022), the APE pipeline lets LLMs generate prompt candidates and then selects from them. In our experiments, for each task, following original templates, $30$ prompts are generated, followed by different methods to perform selection with a budget of $5$ evaluations per prompt on average (i.e., $150$ LLM accesses overall). Zhou et al., (2022) suggest a non-iterative version with uniform evaluations of prompts, which is taken as the baseline here.
@@
The end-to-end experiment results are reported in Table 3, which reveal the consistently better performance of TRIPLE over the originally adopted baseline methods. This observation highlights the applicability and flexibility of the TRIPLE framework, i.e., it can benefit any prompt optimization pipelines requiring a selection component.
6 Extension: Selections of Examples for Few-shot Prompts
Based on the general connection between prompt optimization and BAI-FB, the power of TRIPLE can be further extended beyond finding one good instructional prompt. In the following, we provide discussions on how to leverage TRIPLE to efficiently select examples for few-shot prompts.
As noticed in Brown et al., (2020), LLMs can perform varying tasks when prompted with several related examples, i.e., few-shot prompting. It has been widely recognized that a good choice of examples in few-shot prompts is important to obtain good downstream performances (Liu et al., 2021a, ; Lu et al.,, 2021). Using the terminology introduced in Sec. 1, we can formulate the problem of example selection as follows. From a set of examples ${\mathcal{G}}$, we target at selecting $M$ examples $(g_{1},\cdots,g_{M})$ to form a few-shot prompt, whose performance is measured as $\mu(g_{1},\cdots,g_{M}):=\mathbb{E}_{X\sim{\mathcal{I}}_{X}}\mathbb{E}_{\hat{Y}\sim f([g_{1},\cdots,g_{M};X])}[s(X,\hat{Y})]$. The optimal selection of examples can be defined as $(g^{*}_{1},\cdots,g^{*}_{M}):=\operatorname*{arg\,max}_{g_{1},\cdots,g_{M}\in{\mathcal{G}}}\mu(g_{1},\cdots,g_{M})$.
From the MAB perspective, the learning target can be interpreted as finding the optimal combination of $M$ arms from the overall arm set ${\mathcal{G}}$, which is the focus of the study on combinatorial MAB (CMAB) (Chen et al.,, 2013, 2016; Combes et al.,, 2015). Then, TRIPLE can be further extended to incorporate BAI-FB designs from CMAB to perform the desired example selection. In particular, based on some heuristics on the performance $\mu(g_{1},\cdots,g_{M})$, TRIPLE-SAR and TRIPLE-CSAR are proposed, extending Chen et al., (2014); Gabillon et al., (2011), which are further discussed in Appendix D. The performances of these extensions are presented in Table 4, with more details and results provided in Appendix E.6 and F.
Table 4: Performance comparisons of various example selection methods on different tasks using GPT-3.5 with $|{\mathcal{G}}|=50$ candidate examples, budget $N=100$, and length $M=4$. The tasks are numbered according to Table 3. For each task, the best score across is marked as red, and the second best as yellow.
Tasks Random Uniform SAR CSAR Tasks Random Uniform SAR CSAR
#1 0.65$\pm$ 0.07 0.63$\pm$ 0.13 0.67$\pm$ 0.07 0.66$\pm$ 0.07 #7 0.98$\pm$ 0.03 1.00$\pm$ 0.00 1.00$\pm$ 0.00 1.00$\pm$ 0.00
#2 0.21$\pm$ 0.06 0.26$\pm$ 0.05 0.24$\pm$ 0.07 0.27$\pm$ 0.07 #8 0.35$\pm$ 0.02 0.40$\pm$ 0.05 0.38$\pm$ 0.05 0.42$\pm$ 0.06
#3 0.83$\pm$ 0.06 0.90$\pm$ 0.05 0.93$\pm$ 0.07 0.91$\pm$ 0.06 #9 0.55$\pm$ 0.14 0.64$\pm$ 0.12 0.65$\pm$ 0.12 0.65$\pm$ 0.14
#4 0.96$\pm$ 0.02 0.96$\pm$ 0.01 0.97$\pm$ 0.01 0.97$\pm$ 0.01 #10 0.84$\pm$ 0.10 0.91$\pm$ 0.05 0.95$\pm$ 0.01 0.94$\pm$ 0.05
#5 0.73$\pm$ 0.11 0.80$\pm$ 0.05 0.73$\pm$ 0.05 0.84$\pm$ 0.10 #11 0.41$\pm$ 0.18 0.82$\pm$ 0.20 0.68$\pm$ 0.13 0.87$\pm$ 0.20
#6 0.78$\pm$ 0.10 0.79$\pm$ 0.11 0.84$\pm$ 0.04 0.82$\pm$ 0.04 #12 0.65$\pm$ 0.08 0.56$\pm$ 0.07 0.62$\pm$ 0.12 0.70$\pm$ 0.09
Avg. Performance Rank 3.75$\pm$ 0.59 2.83$\pm$ 0.69 2.08$\pm$ 0.86 1.33$\pm$ 0.47
@@ 7 Conclusions
Prompt optimization is an important problem for large language models (LLMs), but prior research has not considered the potential cost during prompt selection. We have explicitly incorporated a budget constraint to prompt optimization, and studied the problem of how to efficiently select prompts with the given budget. A systematical connection between multi-armed bandits (MAB) and prompt optimization was established. Through this lens, we proposed a general framework, termed TRIPLE, to fully harness the power of fixed-budget best arm identification (BAI-FB) to perform prompt optimization. Besides standard BAI-FB designs, two embedding-based enhancements were proposed to accelerate learning. Extensive experimental results demonstrated the superiority of TRIPLE over multiple representative tasks and various targeted LLMs. Furthermore, we showed that TRIPLE could be plugged into popular end-to-end prompt optimization pipelines, with better performance than previous implementations, demonstrating its effectiveness and flexibility.
@@
In addition to the technical contributions, we believe that the connection between prompt optimization and MAB may be of broader interest. It not only provides a rich set of tools from MAB to advance prompt optimization but also introduces a new application scenario for MAB (especially BAI) research. In particular, the discussed extension to the selection of examples for few-shot prompts demonstrates the rich potential of TRIPLE. As future steps, the research on contextual bandits (Li et al.,, 2010) may provide insights into selecting input-dependent prompts (Wu et al.,, 2022). Also, the application of prompt optimization may spark new research efforts in MAB, e.g., efficient BAI methods for correlated arms.
Appendix A Related Works
Prompt Optimization. The study of prompt optimization (also known as instruction learning) focuses on automatically learning suitable prompts, which is more scalable compared with manual prompt engineering. Many efforts have been devoted to this direction (Gao et al.,, 2020; Wen et al.,, 2023; Sun et al.,, 2023; Zhang et al.,, 2022; Zhang et al., 2023b, ), which is summarized in a recent survey of Liu et al., (2023). The early studies mostly consider optimizing soft prompts (i.e., continuous vectors) (Lester et al.,, 2021; Li and Liang,, 2021; Zhong et al.,, 2021; Liu et al., 2021b, ; Wu et al.,, 2022) or discrete but not interpretable prompts (Shin et al.,, 2020; Shi et al.,, 2022) for white-box LLMs, where different gradient-guided optimization techniques are leveraged. The recent research efforts, instead, are more focused on considering the more practical setting of learning interpretable prompts for black-box LLMs (Prasad et al.,, 2022; Zhou et al.,, 2022; Pryzant et al.,, 2023; Xu et al.,, 2022; Guo et al.,, 2023; Pan et al.,, 2023; Chen et al.,, 2023), where the generating-then-selecting pipeline in Fig. 1 is often adopted.
Compared with previous investigations, this work additionally considers a limited budget in an explicit fashion, which we believe is a practical but under-investigated concern. Most of the previous methods perform selection based on evaluating prompts on all development data (Jiang et al.,, 2020; Xu et al.,, 2022; Guo et al.,, 2023; Prasad et al.,, 2022), which is unavoidably costly. APE (Zhou et al.,, 2022) briefly touched on the cost issue by proposing a naive iterative top filtering strategy (which however is suggested to have only marginal benefits). APO (Pryzant et al.,, 2023) tested a few MAB methods, including two classical BAI-FB designs, i.e., SH (Karnin et al.,, 2013) and SR (Audibert et al.,, 2009), and reported that the performance of UCB is the more favorable. However, it neither formally introduces the budget constraint nor provides a systematical connection to BAI-FB as in this work. Also, this work goes much deeper in incorporating the state-of-the-art BAI-FB designs (i.e., CR (Wang et al.,, 2023) and GSE (Azizi et al.,, 2023)) and proposing embedding-based enhancements. It is worth noting that INSTINCT (Lin et al.,, 2023) also incorporates one MAB method, i.e., NeuralUCB (Zhou et al.,, 2020), to prompt optimization. However, as mentioned in Sec. 1, NeuralUCB is designed for regret minimization (instead of best arm identification), which is not suitable for learning the optimal prompt.
Multi-armed Bandits. Here we briefly discuss the representative studies of MAB, with comprehensive surveys available in Bubeck et al., (2012); Lattimore and Szepesvári, (2020). As mentioned in Sec. 3.1, the target of learning in a MAB system can be roughly categorized as regret minimization and best arm identification. The regret minimization designs target achieving a desired balance between exploration and exploitation so that the cumulative rewards are maximized, e.g., UCB (Auer et al.,, 2002) and Thompson sampling (Thompson,, 1933). The best arm identification (BAI) designs are fully focused on exploration and can be further divided into two classes: fixed-budget (BAI-FB) and fixed-confidence (BAI-FC). The BAI-FB setting maximizes the probability of finding the best arm with a limited number of pulls (Audibert et al.,, 2010; Karnin et al.,, 2013; Wang et al.,, 2023; Azizi et al.,, 2023; Atsidakou et al.,, 2022; Yang and Tan,, 2022). The BAI-FC setting is a dual one which focuses on minimizing the overall number of pulls while guaranteeing that the probability of finding the best arm is higher than a threshold (Garivier and Kaufmann,, 2016; Jamieson and Nowak,, 2014; Soare et al.,, 2014; Jedra and Proutiere,, 2020). Besides leveraging BAI-FB as in this work, it is imaginable that BAI-FC designs can also find applications in prompt optimization, especially when valuing identification accuracy over incurred costs. While MAB has found wide success in different applications, this work marks the first time that a systematical connection between MAB and prompt optimization has been established to the best of our knowledge.
Appendix B Discussions
B.1 Broader Impacts
This work introduces TRIPLE, a framework that can perform efficient prompt optimization for large language models (LLMs) under limited budgets. By optimizing resource usage in prompt optimization for LLMs, we believe the proposed approach could make advanced AI tools and research more accessible to institutions and individuals with limited budgets, promoting a more equitable and democratized field of study. While acknowledging the need for responsible usage of the proposed method, we do not foresee major negative societal impacts.
B.2 Limitations and Future Works
This work opens an interesting direction on the connection between MAB and prompt optimization. In the following, we discuss a few aspects that are currently lacking in this work and particularly worth future explorations.
$\bullet$ Prompt-specific costs. This work considers an abstract model where the cost during learning is measured by the number of LLM accesses. This model provides an important starting point to initialize the investigation. To make the study more practical, more refined considerations on costs can be incorporated. For example, the OpenAI API charge the interactions based on the number of input tokens, which means longer prompts incur higher costs. The cost-aware BAI studied in Kanarios et al., (2024) can provide some insights to further considering prompt-specific costs.
$\bullet$ Other BAI designs. Based on the connection between prompt optimization and BAI, this work has incorporated several BAI designs. However, the research on MAB has a long and rich history, where many other BAI designs can also be leveraged. For example, the Bayesian perspective provided in Komiyama et al., (2023); Atsidakou et al., (2022) and the function approximation scheme adopted in Yavas and Tan, (2023); Yang and Tan, (2022) are all worth investigations. This work is important in delivering the message that (both existing and forthcoming) BAI methods can benefit prompt optimization, which may inspire future explorations.
$\bullet$ Structured prompts. In Sec. 6, we discuss how to extend TRIPLE to select examples for few-shot prompts. Based on the insights obtained in this work, we believe this direction is work further exploration. Moreover, other forms of structured prompting methods, such as Chain-of-Thoughts (Wei et al.,, 2022), are also interesting topics, which may further require multi-step techniques such as reinforcement learning.
Appendix C Details of TRIPLE Designs